OpenStudy (sushi121212):

Factor and simplify: (a^2+2a)^2-2(a^2+2a)-3

OpenStudy (decentnabeel):

\[\left(a^2+2a\right)^2:\quad a^4+4a^3+4a^2\] \[=a^4+4a^3+4a^2-2\left(a^2+2a\right)-3\] \[\mathrm{Expand}\:-2\left(a^2+2a\right):\quad -2a^2-4a\] \[=a^4+4a^3+4a^2-2a^2-4a-3\] \[\mathrm{Add\:similar\:elements:}\:4a^2-2a^2=2a^2\] \[=a^4+4a^3+2a^2-4a-3\]

OpenStudy (sushi121212):

Thank you!

OpenStudy (decentnabeel):

\[\mathrm{Factor}\:a^4+4a^3+2a^2-4a-3:\quad \left(a+1\right)^2\left(a+3\right)\left(a-1\right)\]

OpenStudy (sushi121212):

Is that the answer?

OpenStudy (sushi121212):

Could you tell me how to factor it?

OpenStudy (decentnabeel):

\[\mathrm{Factor}\:a^4+4a^3+2a^2-4a-3:\quad \left(a+1\right)^2\left(a+3\right)\left(a-1\right)\] \[a^4+4a^3+2a^2-4a-3\] Use the rational root theorem \[a=3,\:\quad a_n=1\] \[\mathrm{The\:dividers\:of\:}a:\quad 1,\:3,\:\quad \mathrm{The\:dividers\:of\:}a_n:\quad 1\] \[\mathrm{Therefore,\:\check\:the\:following\:rational\:numbers:\quad }\pm \frac{1,\:3}{1}\] \[-\frac{ 1 }{ 1 }\] is a root of the expression,so factor out a+1 \[\mathrm{Compute\:}\frac{a^4+4a^3+2a^2-4a-3}{a+1}\mathrm{\:\to\:get\:the\:rest\:of\:the\:eqution:\quad }a^3+3a^2-a-3\] \[=\left(a+1\right)\left(a^3+3a^2-a-3\right)\] \[\mathrm{Factor}\:a^3+3a^2-a-3:\quad \left(a+3\right)\left(a+1\right)\left(a-1\right)\] \[=\left(a+1\right)\left(a+3\right)\left(a+1\right)\left(a-1\right)\] \[=\left(a+1\right)^2\left(a+3\right)\left(a-1\right)\]